11.

11. OlZWNzteRaPL5uOBZ9lg9pblLk6BN9m5xWhjNbnC
Answer #1

Answer 11

Net Present Worth = -Initial Cost + Present Value of net annual benefit (H,ere net annual benefit =Additional revenue - Additional labor and maintenance cost = 311750 - 120250 = 191500) + Present Value of salvage Value(Present value when this is sold as scrap)

Present Value(PV) of periodic payment is given by :

PV = (P/r)(1 - 1/(1 + r)n)

Here PV = Present value, P = Annual net benefit = 191500, r = MARR = 12% = 0.12 and n = 10

Hence Present Value of net annual benefit = (191500/0.12)(1 - 1/(1 + 0.12)10) = 1082017.71

Present value of amount received after n years is given by :

PV = A/(1 + r)n

where A = Amount after n years = 24,300, r = MARR = 12% = 0.12 and n = 10

Hence PV = 24300/(1 + 0.12)10 = 7823.95

Hence

Net Present Worth = -Initial Cost + Present Value of net annual benefit (H,ere net annual benefit =Additional revenue - Additional labor and maintenance cost = 311750 - 120250 = 191500) + Present Value of salvage Value(Present value when this is sold as scrap)

=> Net Present Worth = -1,000,000 + 1082017.71 + 7823.95 = 89841.66

Thus, Net Present Worth > 0. Hence, This machine should be purchased.